# Reservoir-Engineered Mechanical Cat States with a Driven Qubit

### v1

@@ III.2 Cat-State Generation
We simulate the mechanical dynamics under the effective master equation (III.1) to track how these features emerge in time. We adopt experimentally accessible parameters Xiang et al. (2013); O’Connell et al. (2010); Niemczyk et al. (2010); Paauw et al. (2009); Chu et al. (2017b): $g_{z}/2\pi=6~\mathrm{MHz}$, $g_{x}=0.1\,g_{z}$; mechanical frequency $\omega_{m}/2\pi=100~\mathrm{MHz}$; qubit decay rate $\kappa/2\pi=100~\mathrm{kHz}$; and mechanical damping $\gamma/2\pi=15~\mathrm{Hz}$. We set the qubit and drive resonantly to $\omega_{q}=\omega_{d}=2\omega_{m}$ to activate two-phonon processes, and we take a zero-temperature mechanical bath in the numerics (i.e., $n_{\mathrm{th}}=0$). The mechanical Hilbert space is truncated to 60 to ensure convergence. Fig. 1 displays snapshots of $W(x,p)$: starting from the vacuum (single Gaussian peak), the distribution splits into two lobes and develops interference fringes at intermediate times, and by $\Gamma_{2}^{-}t\simeq 1$ it approaches a steady even-parity cat state with two well-separated peaks. A detailed comparison between the full master equation [Eq. (2)] and the reduced effective model [Eq. (III.1)] is presented in Appendix C.
@@
In parallel, interference between the coherent drive and the nonlinear mixing generated by the combined couplings produces a coherent two-phonon squeezing term in the reduced Hamiltonian, so that on resonance $\omega_{d}=\omega_{q}$ one has $|\chi|=2\varepsilon g/\kappa$. The coexistence of strong, parity-preserving two-phonon cooling ($\Gamma_{2}^{-}\gg\gamma$) and coherent two-phonon squeezing produces a steady state with robust interference fringes and a characteristic two-lobe structure in the Wigner function. This behavior appears over a broad but specific hierarchy: $g_{x,z}\ll\omega_{m}$ (validity of the Schrieffer–Wolff step), $\kappa\gg\{\Gamma_{2}^{-},\,\Gamma_{1}^{\pm},\,\gamma\}$ (adiabatic elimination), and $\omega_{q}\simeq\omega_{d}\simeq 2\omega_{m}$ (two-phonon resonance). Related second-order, drive-assisted phonon-pair processes have been demonstrated in superconducting-qubit platforms Lambert et al. (2018), consistent with the operating regime considered here.
@@ IV Conclusion
We have introduced a reservoir-engineering scheme for preparing Schrödinger-cat states in a nanomechanical resonator by coupling it to a coherently driven, dissipative two-level system. The core mechanism relies on exploiting both longitudinal and transverse couplings between the qubit and the mechanical mode while driving the qubit near twice the oscillator frequency. This configuration activates resonant two-phonon exchange processes, enabling energy from the drive to be coherently converted into phonon pairs within the resonator. To describe the resulting dynamics, we described the microscopic model based on a time-dependent Hamiltonian and applied a sequence of transformations. First, we used a Schrieffer–Wolff transformation to capture the effective nonlinear interactions. Then, by adiabatically eliminating the strongly dissipative qubit using the Nakajima–Zwanzig formalism, we derived an effective master equation for the mechanical resonator. The reduced equation features (i) qubit-induced single- and two-phonon dissipators with rates determined by qubit response functions, (ii) a coherent two-phonon squeezing term with amplitude $\chi$ set by the drive–qubit detuning and linewidth, and (iii) qubit-mediated frequency shifts together with an effective Kerr nonlinearity. Their combined action drives the oscillator toward a nonclassical steady state characterized by well-defined parity and quantum interference in the Wigner function—hallmarks of a Schrödinger-cat state.
@@ Appendix A Experimental Platform and Feasibility
To implement our scheme for the dissipative generation of mechanical Schrödinger cat states, we require a controllable quantum platform that supports both longitudinal and transverse qubit–resonator interactions. A superconducting charge qubit such as a Cooper-pair box (CPB) or transmon, coupled to a nanomechanical resonator provides such an architecture, with both coupling types realizable through careful circuit design. Coherent control is enabled via external drives, while relaxation and decoherence arise from natural or engineered couplings to electromagnetic and phononic baths. Here, we review and reconstruct the system Hamiltonian and master equation relevant to our proposed scheme. Our goal is to show how the model used in Eq. (2) of the main text can be realized within a well-established superconducting qubit–mechanical resonator architecture Armour et al. (2002); Xiang et al. (2013). Simultaneous longitudinal and transverse coupling between superconducting qubits and mechanical resonators has been proposed theoretically and considered feasible within circuit QED architectures in several previous works (e.g., Niemczyk et al. (2010); Zhang et al. (2018)). For the CPB, the longitudinal and transverse couplings both arise from the same capacitive pathway after rotation to the qubit eigenbasis; no extra circuit element is required beyond choosing the bias point Makhlin et al. (2001); Niemczyk et al. (2010); Blais et al. (2021). Here, for completeness and clarity, we now describe the physical origin of each term in the Hamiltonian and explain the system’s coupling to environmental degrees of freedom. Building on these foundations, we further show how coherent control and dissipation mechanisms can be incorporated in a consistent manner to yield the master equation (2).
@@
This generalized coupling structure, including both longitudinal and transverse interactions, plays a crucial role in enabling nonlinear effects necessary for engineered dissipation and cat-state generation, and has been studied in prior works such as Refs. Niemczyk et al. (2010); Zhang et al. (2018). We next consider a coherent drive applied simultaneously to the gate and the flux. The gate drive modulates the offset charge as $n_{d}(t)=n_{d}\cos(\omega_{d}t)$ Nakamura et al. (1999), while the flux drive (for a split-junction CPB) modulates the Josephson energy as $E_{J}(t)=E_{J}^{(0)}+\delta E_{J}\cos(\omega_{d}t)$ Didier et al. (2019). The resulting drive Hamiltonian in the charge basis is
$H_{d}^{(\text{charge})}(t)=4E_{C}n_{d}\cos\omega_{d}t\,\sigma_{z}^{(\text{charge})}+\frac{\delta E_{J}}{2}\cos\omega_{d}t\,\sigma_{x}^{(\text{charge})}.$ (31)
Transforming to the eigenbasis using $\sigma_{z}^{(\text{charge})}=\sigma_{z}\cos\theta_{0}+\sigma_{x}\sin\theta_{0}$ and $\sigma_{x}^{(\text{charge})}=-\sigma_{z}\sin\theta_{0}+\sigma_{x}\cos\theta_{0}$, we obtain
$H_{d}(t)=\cos\omega_{d}t\,[A_{z}\sigma_{z}+A_{x}\sigma_{x}],\qquad A_{z}=4E_{C}n_{d}\cos\theta_{0}-\frac{\delta E_{J}}{2}\sin\theta_{0},\quad A_{x}=4E_{C}n_{d}\sin\theta_{0}+\frac{\delta E_{J}}{2}\cos\theta_{0}.$ (32)
A purely transverse drive ($A_{z}=0$) is achieved by phase- and amplitude-matching the two drives:
$\delta E_{J}=8E_{C}\,n_{d}\cot\theta_{0}.$ (33)
Under this condition the transverse amplitude becomes $A_{x}=4E_{C}n_{d}/\sin\theta_{0}$, so that the drive Hamiltonian takes the form
$H_{d}(t)=\Omega\cos\omega_{d}t\;\sigma_{x},\qquad\Omega\equiv\frac{4E_{C}n_{d}}{\sin\theta_{0}}.$ (34)
We note that, in practice, the effective $\sigma_{x}$ drive can be realized by phase-locked modulation of the qubit’s gate and/or its flux bias Krantz et al. (2019); Blais et al. (2021); Vion et al. (2002); Makhlin et al. (2001). This is routinely done in circuit QED to implement sideband transitions and tunable/parametric couplers Leek et al. (2009); Strand et al. (2013); Caldwell et al. (2018); Didier et al. (2015) and does not require two independent free-running sources Krantz et al. (2019); Blais et al. (2021). In our scheme, the relative amplitude and phase are chosen so that the residual $\sigma_{z}$ component cancels, yielding a purely transverse drive Makhlin et al. (2001); Blais et al. (2021). This can be implemented with existing CPB control lines Nakamura et al. (1999); Vion et al. (2002), and can even be reduced to a single-port (flux-only) implementation in suitably designed devices using parametric/sideband driving Strand et al. (2013); Leek et al. (2009); Didier et al. (2015). While we work with a CPB here for concreteness, the same Hamiltonian structure and adiabatic-elimination procedure apply to a fluxonium qubit capacitively or piezoelectrically coupled to a nanomechanical mode Manucharyan et al. (2009); Blais et al. (2021); Nguyen et al. (2019). Fluxonium supports MHz–GHz-range transition frequencies with large anharmonicity, which makes the two-phonon resonance condition $\omega_{q}\approx 2\omega_{m}$ accessible for $\omega_{m}\!\sim\!10^{2}\text{--}10^{3}$ MHz. Moreover, dispersive coupling between superconducting qubits and mechanical oscillators has been demonstrated experimentally, and standard Purcell engineering of $\kappa$ is available in circuit QED Blais et al. (2021).
@@ Appendix B Derivation of the Effective Master Equation
In this appendix, we present a detailed derivation of the second-order contributions to the effective master equation governing the reduced dynamics of the mechanical resonator, presented in Eq. (III.1). Our goal is to adiabatically eliminate the qubit degrees of freedom and derive a closed master equation for the resonator, valid in the regime where the qubit rapidly relaxes to its steady state due to strong dissipation. The effective interaction Hamiltonian, obtained via a Schrieffer–Wolff transformation [Eq. (9)], is transformed into the interaction picture with respect to the free evolution of the qubit and the resonator. In this frame, the interaction Hamiltonian decomposes as
@@
$\frac{d}{dt}\mathcal{P}\rho(t)=\mathcal{P}\mathcal{L}\mathcal{P}\rho(t)+\int_{0}^{\infty}ds\,\mathcal{P}\mathcal{L}(t)e^{\mathcal{Q}\mathcal{L}_{0}s}\mathcal{Q}\mathcal{L}(t-s)\mathcal{P}\rho(t),$ (41)
@@
Single-phonon interaction: Derivation of $M_{11}$
@@
The Eq. (51) describes dissipative one-phonon processes induced by the fast-relaxing qubit, with rates determined by the spectral overlap of the drive and the qubit linewidth. The Lamb shift $\delta_{1}$ arises from virtual excitation and de-excitation of the qubit and leads to a frequency renormalization of the mechanical mode.
Two-phonon interaction: Derivation of $M_{22}$
@@
Two-phonon–drive cross term: Derivation of $M_{23}$
@@
$F_{+}(t,\tau)$ $=e^{i(\omega_{d}+\omega_{q})(t-\tau)}+e^{i(\omega_{q}-\omega_{d})(t-\tau)},\qquad F_{-}(t,\tau)=e^{i(\omega_{d}-\omega_{q})(t-\tau)}+e^{-i(\omega_{d}+\omega_{q})(t-\tau)}.$ (65)

### v2

@@ III.2 Cat-State Generation
We simulate the mechanical dynamics under the effective master equation (III.1) to track how these features emerge in time. We adopt experimentally accessible parameters Xiang et al. (2013); O’Connell et al. (2010); Niemczyk et al. (2010); Paauw et al. (2009); Chu et al. (2017b): $g_{z}/2\pi=6~\mathrm{MHz}$, $g_{x}=0.1\,g_{z}$; mechanical frequency $\omega_{m}/2\pi=100~\mathrm{MHz}$; qubit decay rate $\kappa/2\pi=100~\mathrm{kHz}$; and mechanical damping $\gamma/2\pi=15~\mathrm{Hz}$. We set the qubit and drive resonantly to $\omega_{q}=\omega_{d}=2\omega_{m}$ to activate two-phonon processes, and we take a zero-temperature mechanical bath in the numerics (i.e., $n_{\mathrm{th}}=0$). The mechanical Hilbert space is truncated to 60 to ensure convergence. Fig. 1 displays snapshots of $W(x,p)$: starting from the vacuum (single Gaussian peak), the distribution splits into two lobes and develops interference fringes at intermediate times, and by $\Gamma_{2}^{-}t\simeq 1$ it approaches a steady even-parity cat state with two well-separated peaks. A detailed comparison between the full master equation [Eq. (2)] and the reduced effective model [Eq. (III.1)] is presented in Appendix C. An approximate analytical expression for the steady–state mean phonon number in the cat–state regime, based on a two–phonon cooling picture, is derived in Appendix E.
@@
In parallel, interference between the coherent drive and the nonlinear mixing generated by the combined couplings produces a coherent two-phonon squeezing term in the reduced Hamiltonian, so that on resonance $\omega_{d}=\omega_{q}$ one has $|\chi|=2\varepsilon g/\kappa$. The coexistence of strong, parity-preserving two-phonon cooling ($\Gamma_{2}^{-}\gg\gamma$) and coherent two-phonon squeezing produces a steady state with robust interference fringes and a characteristic two-lobe structure in the Wigner function. This behavior appears over a broad but specific hierarchy: $g_{x,z}\ll\omega_{m}$ (validity of the polaron-like transformation given step in Eq. (7)), $\kappa\gg\{\Gamma_{2}^{-},\,\Gamma_{1}^{\pm},\,\gamma\}$ (adiabatic elimination), and $\omega_{q}\simeq\omega_{d}\simeq 2\omega_{m}$ (two-phonon resonance). Related second-order, drive-assisted phonon-pair processes have been demonstrated in superconducting-qubit platforms Lambert et al. (2018), consistent with the operating regime considered here. The impact of a finite–temperature phonon bath, the induced Kerr nonlinearity, and the degree of sideband resolution on the cat state dynamics is analyzed in detail in Appendix D.
@@ IV Conclusion
We have introduced a reservoir-engineering scheme for preparing Schrödinger-cat states in a nanomechanical resonator by coupling it to a coherently driven, dissipative two-level system. The core mechanism relies on exploiting both longitudinal and transverse couplings between the qubit and the mechanical mode while driving the qubit near twice the oscillator frequency. This configuration activates resonant two-phonon exchange processes, enabling energy from the drive to be coherently converted into phonon pairs within the resonator. To describe the resulting dynamics, we started from a microscopic time-dependent Hamiltonian and applied a sequence of transformations. First, we used a unitary transformation to capture the effective nonlinear interactions. Then, by adiabatically eliminating the strongly dissipative qubit using the Nakajima–Zwanzig formalism, we derived an effective master equation for the mechanical resonator. The reduced equation features (i) qubit-induced single- and two-phonon dissipators with rates determined by qubit response functions, (ii) a coherent two-phonon squeezing term with amplitude $\chi$ set by the drive–qubit detuning and linewidth, and (iii) qubit-mediated frequency shifts together with an effective Kerr nonlinearity. Their combined action drives the oscillator toward a nonclassical steady state characterized by well-defined parity and quantum interference in the Wigner function—hallmarks of a Schrödinger-cat state.
@@ Appendix A Experimental Platform and Feasibility
To implement our scheme for the dissipative generation of mechanical Schrödinger-cat states, we require a controllable quantum platform that supports both longitudinal and transverse qubit–resonator interactions. A superconducting charge qubit such as a Cooper-pair box (CPB) or transmon or fluxonium, coupled to a nanomechanical resonator provides such an architecture, with both coupling types realizable through careful circuit design. Coherent control is enabled via external drives, while relaxation and decoherence arise from natural or engineered couplings to electromagnetic and phononic baths. Here, we review and reconstruct the system Hamiltonian and master equation relevant to our proposed scheme. Our goal is to show how the model used in Eq. (2) of the main text can be realized within a well-established superconducting qubit–mechanical resonator architecture Armour et al. (2002); Xiang et al. (2013). Simultaneous longitudinal and transverse coupling between superconducting qubits and mechanical resonators has been proposed theoretically and considered feasible within circuit QED architectures in several previous works (e.g., Niemczyk et al. (2010); Zhang et al. (2018)). For the CPB, the longitudinal and transverse couplings both arise from the same capacitive pathway after rotation to the qubit eigenbasis; no extra circuit element is required beyond choosing the bias point Makhlin et al. (2001); Niemczyk et al. (2010); Zhang et al. (2018); Ma et al. (2019); Hou et al. (2024); Blais et al. (2021). Here, for completeness and clarity, we now describe the physical origin of each term in the Hamiltonian and explain the system’s coupling to environmental degrees of freedom. Building on these foundations, we further show how dissipation mechanisms can be incorporated in a consistent manner to yield the master equation (2).
@@
This generalized coupling structure, including both longitudinal and transverse interactions, plays a crucial role in enabling nonlinear effects necessary for engineered dissipation and cat-state generation Ma et al. (2019); Hou et al. (2024); Hauer et al. (2023). The coherent drive term introduced in Eq. (1) can be implemented experimentally using an external microwave field, as discussed in prior studies Ma et al. (2019); Hou et al. (2024); Radić2022; Guthrie et al. (2022). While we work with a CPB here for concreteness, the same Hamiltonian structure and adiabatic-elimination procedure apply to a fluxonium qubit capacitively or piezoelectrically coupled to a nanomechanical mode Manucharyan et al. (2009); Blais et al. (2021); Nguyen et al. (2019). Fluxonium supports MHz–GHz-range transition frequencies with large anharmonicity, which makes the two-phonon resonance condition $\omega_{q}\approx 2\omega_{m}$ accessible for $\omega_{m}\!\sim\!10^{2}\text{--}10^{3}$MHz. Moreover, dispersive coupling between superconducting qubits and mechanical oscillators has been demonstrated experimentally, and standard Purcell engineering of $\kappa$ is available in circuit QED Blais et al. (2021).
@@ Appendix B Derivation of the Effective Master Equation
In this appendix, we present a detailed derivation of the second-order contributions to the effective master equation governing the reduced dynamics of the mechanical resonator, presented in Eq. (III.1). Our goal is to adiabatically eliminate the qubit degrees of freedom and derive a closed master equation for the resonator, valid in the regime where the qubit rapidly relaxes to its steady state due to strong dissipation. The effective interaction Hamiltonian, obtained via a unitary transformation [Eq. (9)], is transformed into the interaction picture with respect to the free evolution of the qubit and the resonator. In this frame, the interaction Hamiltonian decomposes as
@@
$\frac{d}{dt}\mathcal{P}\rho(t)=\mathcal{P}\mathcal{L}\mathcal{P}\rho(t)+\int_{0}^{\infty}\!d\tau\,\mathcal{P}\mathcal{L}(t)e^{\mathcal{Q}\mathcal{L}_{0}\tau}\mathcal{Q}\mathcal{L}(t-s)\mathcal{P}\rho(t),$ (37)
@@
B.1 Single-phonon interaction: Derivation of $M_{11}$
@@
Eq. (47) describes dissipative one-phonon processes induced by the fast-relaxing qubit, with rates determined by the spectral overlap of the drive and the qubit linewidth. The Lamb shift $\delta_{1}$ arises from virtual excitation and de-excitation of the qubit and leads to a frequency renormalization of the mechanical mode.
B.2 Two-phonon interaction: Derivation of $M_{22}$
@@
B.3 Two-phonon–drive cross term: Derivation of $M_{23}$
@@
$F_{\pm}(t,\tau)=e^{\pm i(\omega_{q}+\omega_{d})(t-\tau)}+e^{\pm i(\omega_{q}-\omega_{d})(t-\tau)}.$ (61)
@@
B.4 Why Only Three Second-Order Kernels Survive
To make the selection of second–order contributions fully transparent, this subsection provides explicit calculations showing which Nakajima–Zwanzig kernels in Eq. (38) vanish and why. Starting from the nine kernels $M_{ij}$ generated by $H_{1}$, $H_{2}$, and $H_{3}$, and building on the explicit evaluations of $M_{11}$, $M_{22}$, and $M_{23}$ in the preceding subsections, we demonstrate that only these three kernels survive. The remaining six fall into three natural groups, each eliminated by the same mechanism. Group I (outer drive commutator), $M_{31}$, $M_{32}$, and $M_{33}$, vanishes because the partial trace of a qubit–only commutator is identically zero. Group II (mixed one– and two–phonon vertices), $M_{12}$ and $M_{21}$, reduces to cubic mechanical monomials that carry only fast phases at $\pm\omega_{m}$ or $\pm 3\omega_{m}$ and therefore average out under the RWA. Group III (one–phonon and drive), $M_{13}$, is off–resonant; with $\omega_{d}\simeq 2\omega_{m}$, its residual detunings are $\omega_{d}\pm\omega_{m}$, so no near–stationary contribution remains after coarse–graining. Throughout, the eliminations follow two explicit criteria: (i) a rotating–wave rule, whereby only near–zero net frequencies are retained, and (ii) a mechanical selection rule, whereby odd mechanical monomials do not yield stationary generators in the coarse–grained dynamics.
B.4.1 Group I: Vanishing kernels with an outer drive commutator ($M_{31}$, $M_{32}$, $M_{33}$)
Here we show that all kernels whose outer commutator involves the drive Hamiltonian $H_{3}(t)$ vanish:
$M_{31}(t)\rho_{m}=0,\qquad M_{32}(t)\rho_{m}=0,\qquad M_{33}(t)\rho_{m}=0.$ (74)
The cancellations originate from the fact that, after expanding the inner commutator in Eq. (38) and propagating with $e^{\mathcal{L}_{0}\tau}$, the outer commutator with $H_{3}(t)$ produces a factor of $\sigma_{z}$ acting only on the qubit, whose partial trace is zero. In what follows, we repeatedly use the cyclicity identity under a partial trace: for any qubit operator $A_{q}$ and any bipartite $X=\sum_{k}O_{q,k}\otimes O_{m,k}$,
$\operatorname{Tr}_{q}\big([A_{q},X]\big)=0,$ (75)
and the elementary property $\operatorname{Tr}_{q}(\sigma_{z})=0$.
(i) Kernel $M_{31}$: $H_{3}$ outside, $H_{1}$ inside. To evaluate $M_{31}(t)$, we start from Eq. (38) with $i=3$ and $j=1$, and take the qubit steady state as $\rho_{q}^{\mathrm{ss}}=|g\rangle\langle g|$. This gives
$M_{31}(t)\rho_{m}=-\operatorname{Tr}_{q}\!\left[H_{3}(t),\int_{0}^{\infty}d\tau\ e^{\mathcal{L}_{0}\tau}\,\big[H_{1}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}\big]\right].$ (76)
For the inner commutator, we substitute $H_{1}(t)$ from Eq. (41), this yields
$C_{31}(\tau)=\big[H_{1}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}\big]=g_{x}\Big(\sigma_{+}e^{i\omega_{q}(t-\tau)}\otimes B_{1}(t-\tau)\rho_{m}-\sigma_{-}e^{-i\omega_{q}(t-\tau)}\otimes\rho_{m}B_{1}(t-\tau)\Big).$ (77)
Next, we propagate $C_{31}(\tau)$ under the Liouvillian $\mathcal{L}_{0}$, this gives
$e^{\mathcal{L}_{0}\tau}C_{31}(\tau)=g_{x}e^{-\kappa\tau/2}\Big(\sigma_{+}e^{i\omega_{q}(t-\tau)}\otimes B_{1}(t-\tau)\rho_{m}-\sigma_{-}e^{-i\omega_{q}(t-\tau)}\otimes\rho_{m}B_{1}(t-\tau)\Big).$ (78)
We now evaluate the outer commutator in Eq. (76) with the drive Hamiltonian $H_{3}(t)$ given in Eq. (35). Using
$H_{3}(t),\sigma_{+}]=-\,\varepsilon E_{d}(t)\,e^{-i\omega_{q}t}\sigma_{z},\qquad[H_{3}(t),\sigma_{-}]=\,\varepsilon E_{d}(t)\,e^{+i\omega_{q}t}\sigma_{z},$ (79)
we obtain
$\big[H_{3}(t),e^{\mathcal{L}_{0}\tau}C_{31}(\tau)\big]=-\,g_{x}\varepsilon\,E_{d}(t)\,e^{-\kappa\tau/2}\Big(e^{-i\omega_{q}\tau}\,\sigma_{z}\otimes B_{1}(t-\tau)\rho_{m}+e^{+i\omega_{q}\tau}\,\sigma_{z}\otimes\rho_{m}B_{1}(t-\tau)\Big),$ (80)
here $E_{d}(t)=e^{i\omega_{d}t}+e^{-i\omega_{d}t}$. Taking the partial trace over the qubit in Eq. (80) and using
$\operatorname{Tr}_{q}(\sigma_{z}\otimes X_{m})=\operatorname{Tr}_{q}(\sigma_{z})\,X_{m}=0,$ (81)
we immediately obtain
$M_{31}(t)\rho_{m}=0.\,$ (82)
(ii) Kernel $M_{32}$: $H_{3}$ outside, $H_{2}$ inside. Analogously,
$M_{32}(t)\rho_{m}=-\operatorname{Tr}_{q}\!\left[H_{3}(t),\int_{0}^{\infty}d\tau\ e^{\mathcal{L}_{0}\tau}\,\big[H_{2}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}\big]\right],$ (83)
with $H_{2}(t)$ given in Eq. (57). Following the similar procedure as in $M_{31}$, the inner commutator becomes
$C_{32}(\tau)=\big[H_{2}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}\big]=g\Big(\sigma_{+}e^{i\omega_{q}(t-\tau)}\otimes B_{2}(t-\tau)\rho_{m}+\sigma_{-}e^{-i\omega_{q}(t-\tau)}\otimes\rho_{m}B_{2}(t-\tau)\Big).$ (84)
Applying $e^{\mathcal{L}_{0}\tau}$ and Eq. (79), the outer commutator in Eq. (83) modifies to
$\big[H_{3}(t),e^{\mathcal{L}_{0}\tau}C_{32}(\tau)\big]=-\,g\,\varepsilon\,E_{d}(t)\,e^{-\kappa\tau/2}\Big(e^{-i\omega_{q}\tau}\,\sigma_{z}\otimes B_{2}(t-\tau)\rho_{m}-e^{+i\omega_{q}\tau}\,\sigma_{z}\otimes\rho_{m}B_{2}(t-\tau)\Big).$ (85)
Taking the partial trace over the qubit and using the property given in Eq. (81), we obtain
$M_{32}(t)\rho_{m}=0.\,$ (86)
(iii) Kernel $M_{33}$: $H_{3}$ outside, $H_{3}$ inside. Finally, we consider the kernel with the drive Hamiltonian acting on both sides. Starting from
$M_{33}(t)\rho_{m}=-\operatorname{Tr}_{q}\!\left[H_{3}(t),\int_{0}^{\infty}d\tau\ e^{\mathcal{L}_{0}\tau}\,\big[H_{3}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}\big]\right],$ (87)
and using $H_{3}(t)$ given in Eq. (35), the inner commutator followed by propagation under $e^{\mathcal{L}_{0}\tau}$ gives
$e^{\mathcal{L}_{0}\tau}\big[H_{3}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}\big]=\varepsilon\,e^{-\kappa\tau/2}\Big(\sigma_{+}F_{+}(t,\tau)-\sigma_{-}F_{-}(t,\tau)\Big)\otimes\rho_{m},$ (88)
with $F_{\pm}(t,\tau)$ is given in Eq. (61). Taking the outer commutator in Eq. (87) and using Eq. (79), we obtain
$\big[H_{3}(t),\,e^{\mathcal{L}_{0}\tau}[H_{3}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}]\big]=-\,2\,\varepsilon^{2}\,E_{d}(t)\,E_{d}(t-\tau)\,e^{-\kappa\tau/2}\cos(\omega_{q}\tau)\,\sigma_{z}\otimes\rho_{m}.$ (89)
Finally, taking the partial trace over the qubit and invoking Eq. (81), we arrive at
$M_{33}(t)\rho_{m}=0.$ (90)
B.4.2 Group II: Mixed one– and two–phonon vertices ($M_{12}$, $M_{21}$)
The mixed kernels $M_{12}$ and $M_{21}$, which couple the single-phonon and two-phonon vertices, vanish after RWA. Physically, every qubit contraction produces cubic mechanical monomials $\{a^{\dagger 2}a,\;a^{2}a^{\dagger},\;a^{3},\;a^{\dagger 3}\}$ that carry net oscillation phases at $\pm\omega_{m}$ or $\pm 3\omega_{m}$. These fast phases remove any time-independent contribution, so no stationary Lindblad or coherent term survives in the RWA. Below we show this explicitly for both orderings, using Eqs. (38), (41), and (49).
Kernel $M_{12}$: outer $H_{1}$, inner $H_{2}$.
We start from Eq. (38) with $i=1$ and $j=2$, which gives
$M_{12}(t)\rho_{m}=-\operatorname{Tr}_{q}\!\left[H_{1}(t),\;\int_{0}^{\infty}d\tau\,e^{\mathcal{L}_{0}\tau}\,\big[H_{2}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}\big]\right].$ (91)
Using $H_{2}(t)$ given in Eq. (49), the inner commutator becomes
$\big[H_{2}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}\big]=g\left(\sigma_{+}e^{i\omega_{q}(t-\tau)}\!\otimes\!B_{2}(t-\tau)\rho_{m}+\sigma_{-}e^{-i\omega_{q}(t-\tau)}\!\otimes\!\rho_{m}B_{2}(t-\tau)\right).$ (92)
Propagating the qubit operators under $e^{\mathcal{L}_{0}\tau}$ and then forming the outer commutator with $H_{1}(t)$ yields, after tracing out the qubit,
$M_{12}(t)\rho_{m}=g\,g_{x}\!\int_{0}^{\infty}\!d\tau\,e^{-\kappa\tau/2}\left(e^{-i\omega_{q}\tau}\,[B_{1}(t),\,B_{2}(t-\tau)\rho_{m}]+e^{+i\omega_{q}\tau}\,[B_{1}(t),\,\rho_{m}B_{2}(t-\tau)]\right).$ (93)
Every product in Eq. (93) is cubic in the mechanical operators and carries global phases $e^{\pm i\omega_{m}t}$ or $e^{\pm 3i\omega_{m}t}$, together with $\tau$–dependent factors $e^{-i(\omega_{q}\pm 2\omega_{m})\tau}$. The latter define the two–phonon response functions $S_{2\pm}$ in Eq. (53). Collecting terms yields a linear combination of $a^{\dagger 2}a,\rho_{m}]$, $a^{2}a^{\dagger},\rho_{m}]$, $a^{3},\rho_{m}]$, and $a^{\dagger 3},\rho_{m}]$, each multiplied by $e^{\pm i\omega_{m}t}$ or $e^{\pm 3i\omega_{m}t}$. By the secular test, these rapidly oscillating contributions average to zero under coarse–graining, so no time–independent generator remains:
$M_{12}(t)\rho_{m}=0.$ (94)
Kernel $M_{21}$: outer $H_{2}$, inner $H_{1}$.
Interchanging the roles of $H_{1}$ and $H_{2}$ we start from
$M_{21}(t)\rho_{m}=-\operatorname{Tr}_{q}\!\left[H_{2}(t),\;\int_{0}^{\infty}d\tau\,e^{\mathcal{L}_{0}\tau}\,\big[H_{1}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}\big]\right].$ (95)
The inner commutator now reads
$\big[H_{1}(t-\tau),\,|g\rangle\langle g|\otimes\rho_{m}\big]=g_{x}\Big(\sigma_{+}e^{i\omega_{q}(t-\tau)}\!\otimes\!B_{1}(t-\tau)\rho_{m}-\sigma_{-}e^{-i\omega_{q}(t-\tau)}\!\otimes\!\rho_{m}B_{1}(t-\tau)\Big),$ (96)
now propagating under $e^{\mathcal{L}_{0}\tau}$, evaluating the outer commutator with $H_{2}(t)$ defined in Eq. (57), and tracing over the qubit finally yield
$M_{21}(t)\rho_{m}=g\,g_{x}\!\int_{0}^{\infty}\!d\tau\,e^{-\kappa\tau/2}\left(e^{-i\omega_{q}\tau}\,[B_{2}(t),\,B_{1}(t-\tau)\rho_{m}]+e^{+i\omega_{q}\tau}\,[B_{2}(t),\,\rho_{m}B_{1}(t-\tau)]\right).$ (97)
Here $B_{1}(t-\tau)$ contributes $\tau$–phases $e^{\mp i\omega_{m}\tau}$, so the response functions are $S_{\pm}$ of Eq. (45). As in the previous case, every mechanical product is cubic and carries global factors $e^{\pm i\omega_{m}t}$ or $e^{\pm 3i\omega_{m}t}$. By the same stationarity criterion, the coarse–grained kernel vanishes:
$M_{21}(t)\rho_{m}=0.$ (98)
In summary, both mixed one– and two–phonon kernels $M_{12}$ and $M_{21}$ reduce exclusively to fast, cubic mechanical sectors and therefore contribute no stationary term to the effective generator.
B.4.3 Group III: One–phonon vertice and drive ($M_{13}$)